OpenStudy (anonymous):

how long does it take for a tennis ball to drop 1 inch?

OpenStudy (anonymous):

use straight line motion formula, with a=g, v_0=0 I assume, y_0=1/12 ft. solve for y=0

OpenStudy (anonymous):

I don't understand, please tell me the answer, not the equation. thank you though.

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

does anyone know the answer, i can't find it online

OpenStudy (anonymous):

y=y_0+v_0*t+(a*t^2)/2 y_0=1/12 ft v_0=0 ft/s a=g=-32.2 ft/s^2 0=1/12+0*t+(-32.2*t^2)/2 ft (32.2*t^2)/2=1/12 ft solve for t.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

what is the mark between the 32.2 and the t

OpenStudy (anonymous):

i cant see it well

OpenStudy (anonymous):

i lost my glasses

OpenStudy (anonymous):

sorry i asked the wrong question

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

You want to take a look at the acceleration of the ball. the write a proportion of a sort. in/sec=in/sec

OpenStudy (stormfire1):

For a falling object \[h=V_0+\frac{1}{2} g t^2\]Since the initial velocity is 0, this reduces to: \[h=\frac{1}{2} g t^2\]Acceleleration due to gravity is:\[9.8 m/s^2=32.2ft/s^2=386.4in/s^2\]Now plug in the values and solve for t: \[1in=\frac{1}{2} (386.4in/s^2)t^2\]\[t=0.0719s\]